Initially there were 7 liters of blue paint and 10 liters of crimson paint. A paint job uses 20% of the blue and 80% of the crimson paint.
(a) What percentage of the total combined amount of paint is used during the job?
(b) What percentage of the total combined amount of paint remains after the job is finished?

Answer by stanbon(75887)

Initially there were 7 liters of blue paint and 10 liters of crimson paint.
A paint job uses 20% of the blue and 80% of the crimson paint.
(a) What percentage of the total combined amount of paint is used during the job?
Amount Used = 0.20*7 + 0.80*10 = 1.4 + 8 = 9.4 liters
Percent Used: 9.4/(7+10) = 0.5529 = 55.29%
==================================================
(b) What percentage of the total combined amount of paint remains after the job is finished?
Ans: 100% - 55.29% = 44.71%
===================================
Cheers,
Stan H.